The following questions relate to Kyle Company, which manufactures products KA, KB, and KC from a joint process. Joint product costs were $197,000. Additional information follows: If Processed Further Product Units Produced Sales Value at Split-Off Sales Values Additional Costs KA 83,000 $ 272,000 $ 320,000 $ 53,000 KB 58,100 238,000 260,000 43,000 KC 24,900 170,000 220,000 31,000

If Processed Further
Product Units Produced Sales Value at Split-Off Sales Values Additional Costs
KA 83,000 $ 272,000 $ 320,000 $ 53,000
KB 58,100 238,000 260,000 43,000
KC 24,900 170,000 220,000 31,000

Required:

a. Assuming that joint product costs are allocated using the physical quantities (units produced) method, what was the total cost of product KA (including $53,000 if processed further)? (Do not round your intermediate calculations.)

b. Assuming that joint product costs are allocated using the sales value at split-off (net realizable value method), what was the total cost of product KB (including the $43,000 if processed further)? (Do not round your intermediate calculations.)
Answer #1

a) Total Cost of product KA =(197000*83000/(83000+58100+24900))+53000 =(197000*83000/166000) +53000 =98500+53000 = $151500. b